OpenStudy (tabithax3):

The floor area of a home is 827 square feet. Find this area in square nanometers.

OpenStudy (nincompoop):

do you the conversion rate from ft to meters?

OpenStudy (tabithax3):

@nincompoop my professor wrote that 3.28 ft = 1 m

OpenStudy (tabithax3):

and nano = \[10^{-9}\]

OpenStudy (tabithax3):

@mayankdevnani i don't know how to convert ft to nanometer

OpenStudy (mayankdevnani):

\[\large \bf 3.28ft=1 m\] \[\large \bf 1 ft=\frac{1}{3.28}m\] we have to find,827 ft, \[\large \bf 827ft=827 \times \frac{1}{3.28}m=252.13m\] but we know that \(\large \bf 1m=10^{-9}nm\) so,252.13m, \[\large \bf 252.13m=252.13 \times 10^{-9}nm=\color{red}{Answer}\]

OpenStudy (mayankdevnani):

understood? @tabithax3

OpenStudy (tabithax3):

@mayankdevnani the problem i don't get is that the question says \[ft^{2}\] and \[nm^{2}\].

OpenStudy (mayankdevnani):

this doesn't matters

OpenStudy (tabithax3):

@mayankdevnani oh okay, i thought it did. thanks!

OpenStudy (mayankdevnani):

welcome :)